# Learning to Instruct for Visual Instruction Tuning

### v1

@@ Abstract
We propose LIT, an advancement of visual instruction tuning (VIT). While VIT equips Multimodal LLMs (MLLMs) with promising multimodal capabilities, the current design choices for VIT often result in overfitting and shortcut learning, potentially degrading performance. This gap arises from an overemphasis on instruction-following abilities, while neglecting the proactive understanding of visual information. Inspired by this, LIT adopts a simple yet effective approach by incorporating the loss function into both the instruction and response sequences. It seamlessly expands the training data, and regularizes the MLLMs from overly relying on language priors. Based on this merit, LIT achieves a significant relative improvement of up to 9% on comprehensive multimodal benchmarks, requiring no additional training data and incurring negligible computational overhead. Surprisingly, LIT attains exceptional fundamental visual capabilities, yielding up to an 18% improvement in captioning performance, while simultaneously alleviating hallucination in MLLMs.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Performance comparison on a broad range of 16 tasks between LIT and VIT using different models, including TinyLLaVA Qwen2-0.5B (Zhou et al., 2024), LLaVA-1.5 Vicuna-7B (Liu et al., 2024a), and LLaVA-1.5 Vicuna-13B (Liu et al., 2024a). The pretraining phase uses the LLaVA-pretrain-558k dataset, while the fine-tuning phase employs the LLaVA-mix-665k dataset. Clearly, our LIT demonstrates significant improvements in overall performance compared to VIT across different models.
Typically, visual instruction tuning consists of two stages: pre-training and fine-tuning (Liu et al., 2023a; Zhu et al., 2024). In the pre-training stage, a connector is trained to align visual and language data. The fine-tuning phase is similar to instruction tuning in language models, where the model is trained to generate responses based on multimodal instructions. However, recent research has shown that instruction tuning can lead to a series of issues, such as the knowledge degradation (Ghosh et al., 2024) and hallucinations (Rawte et al., 2023; Tonmoy et al., 2024). Similar problems have also been observed in models that employ visual instruction tuning (Huang et al., 2023; Leng et al., 2024). The potential causes may lie in overfitting and shortcut learning during instruction tuning (Sun et al., 2024b). Figure 2 illustrates an example of a shortcut, where the model might ignore visual content and generate responses based solely on language priors. Current advancements (Liu et al., 2024a, b; Bai et al., 2023b; Tong et al., 2024) in MLLMs implicitly mitigate such issues by leveraging larger, higher-quality, and more diverse datasets, larger models, and improved pretrained initializations for visual and language backbones.
In this paper, we propose an orthogonal solution to improving visual instruction tuning: learning to instruct images as a regularizer, which we refer to as Learning to InstrucT (LIT). Specifically, in addition to learning to generate responses to instructions as usual, our method also learn to generate instructions for images that exclude templates, which refer to special tokens and high-frequency, low-information template tokens in the instructions. Our method enhances VIT by: (1) expanding the training content to mitigate overfitting without explicitly enlarging the training set (Section 3.4); and (2) learning to instruct images, which forces the model to focus more on the visual content and prevents learning some shortcuts (Section 2).
We conduct extensive experiments across a range of 16 tasks, comparing our LIT and VIT on different models. As shown in Figure 1, LIT achieves a significant relative improvement of up to 6% over VIT in overall multimodal task performance. Moreover, LIT shows significant potential in alleviating hallucination issues in MLLMs, as demonstrated across four diverse hallucination benchmarks (Section 3.3). Through comprehensive and detailed experiments, LIT demonstrates the following advantages: (1) It improves the ability across different multimodal tasks compared to the conventional visual instruction tuning, especially those focusing more on visual content, such as OCR, captioning and hallucination mitigation. (2) Our method is orthogonal to existing research advancements and can further enhance MLLMs’ performance through simple integration. (3) It shows advantages in improving performance while making minimal compromises in training and inference costs.
@@ 2 Method
In this section, we begin with the problem definition, briefly review Visual Instruction Tuning, and then introduce our proposed method LIT.
Problem Formulation. Taking LLaVA as an example, a typical model architecture comprises a pre-trained large language model $f(\cdot)$ with parameters $\theta_{f}$ with the corresponding tokenizer and embedding layer $t(\cdot)$ (e.g., Vicuna (Peng et al., 2023)), a pre-trained visual feature extractor $g(\cdot)$ with parameters $\theta_{g}$ (e.g., CLIP-ViT-L/14 (Radford et al., 2021)), and a cross-modal connector $h(\cdot)$ with parameters $\theta_{h}$, such as a linear layer or MLP. Let $\theta=\{\theta_{f},\theta_{g},\theta_{h}\}$ denote the set of all parameters. For an image $\mathbf{X}_{V}$ and a related text instruction $\mathbf{X}_{I}$, we obtain the corresponding textual response through the following forward process. As shown in Figure 3, the image $\mathbf{X}_{V}$ is forwarded through the visual feature extractor $g$ and cross-modal connector $h$, mapping it to visual tokens $\mathbf{H}_{V}=h(g(\mathbf{X}_{V}))$ in the language embedding space. $\mathbf{H}_{V}$ is then combined with the language tokens $\mathbf{H}_{I}=t(\mathbf{X}_{I})$ to form a sequence, which is forwarded into LLM $f$ for autoregressive generation of the response sequence $\mathbf{X}_{A}$. The goal of visual instruction tuning is to train the model to exhibit strong multi-modal instruction-following capabilities.
Figure 3: Illustration of the model architecture using LLaVA as an example, along with the data flow for generating responses from an image and instruction.
@@
$\mathcal{L}$ $=-\log p_{\theta}(\mathbf{X}_{A}|\mathbf{X}_{V},\mathbf{X}_{I})$ (1)
$=-\sum_{i=1}^{L_{A}}\log p_{\theta}(\mathbf{X}_{A,i}|\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A,<i}).$
@@
Figure 4: Illustration of the proposed LIT. In addition to learning to generate responses like VIT, LIT also learns to generate instructions that exclude templates.
Figure 5: Cases, each including an image, an instruction, and the responses generated by the VIT and LIT models, respectively. The models are based on LLaVA-1.5 Vicuna-7B. (Left) The VIT model shows better OCR capabilities. (Middle) The VIT model demonstrates robustness and effectively avoids being influenced by misleading information in the language instruction. (Right) The VIT model provides a more comprehensive and accurate image description.
@@
Figure 6: The visual contribution ($\mathrm{VC}$) distributions of VIT and LIT on the VQAv2 training data and DocVQA test data. Experiments are based on TinyLLaVA Qwen2-0.5B.
Learning to Instruct (LIT). Building upon VIT, we propose extending its paradigm through learning to instruct images. Specifically, in addition to learning how to generate a response given an image and an instruction, we also learn how to generate a meaningful instruction for a given image. For a training sample triplet $(\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A})$, we define the loss function $\mathcal{L}$ as the negative log-likelihood of both the instruction and the response conditioned on the image:
$\mathcal{L}$ $=-\log p_{\theta}(\mathbf{X}_{I},\mathbf{X}_{A}|\mathbf{X}_{V})=\underbrace{-\sum_{i=1}^{L_{I}}\log p_{\theta}(\mathbf{X}_{I,i}|\mathbf{X}_{V},\mathbf{X}_{I,<i})}_{\text{Learn to Instruct}}$ (2)
$\underbrace{-\sum_{i=1}^{L_{A}}\log p_{\theta}(\mathbf{X}_{A,i}|\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A,<i})}_{\text{Learn to Respond}}.$
By learning to generate appropriate instructions for images, LIT achieves two key benefits. First, it naturally expands the data the model learns to fit, helping to alleviate potential overfitting to some extent. Second, learning to instruct images ensures that the model focuses on the image content, effectively preventing it from ignoring the visual input and relying solely on language priors to generate responses.
Template Removal. To ensure the model learns meaningful content related to the image, we exclude the learning of certain irrelevant parts of instructions. Such irrelevant context primarily arises from two sources: (1) system templates and (2) task templates. Specifically, system templates refer to the tokens used to guide the MLLMs in adopting the role of a helpful and polite AI assistant, or as the conversational clues distinguishing whether the content is generated by the “USER” or the “ASSISTANT”. System templates can be easily removed when it is added. Task templates refer to tokens that indicate the task type and output format. We identify task templates by calculating the frequency of all sentences in the entire training dataset and selecting the most frequent ones. The detailed removed task templates can be referred to Appendix A. Note that the training data in the pretraining stage is constructed using image-caption pairs, with a set of template prompts serving as instructions, which are unrelated to the image content. Therefore, we apply our method LIT only during the end-to-end fine-tuning stage. In Figure 4, we present an illustration of LIT.
Analysis. To assess the rationality of LIT in alleviating shortcut learning in MLLMs, we conduct experiments to gain deeper insight into which aspects of visual or instructional content are most influential in response prediction. Specifically, we first introduce visual contribution ($\mathrm{VC}$), which quantifies the difference in the log-likelihood of the response when conditioned on both the image and instruction, versus when conditioned solely on the instruction:
@@
where $\mathbf{X}_{V}=\emptyset$ in the implementation denotes that the original visual input is replaced with random noise. In this spirit, $\mathrm{VC}$ captures the relative importance of visual signals in response prediction, isolating the contribution of visual information from any confounding effects due to the instructional signals. Figure 6 presents the VC of LIT and VIT for both the training and testing datasets. Our findings reveal that LIT achieves a substantial relative improvement, with a 9% increase in VC, highlighting its effectiveness in regularizing MLLMs to better utilize visual inputs.
Moreover, Figure 5 presents several cases, each consisting of an image, an instruction, and the responses generated by the VIT and LIT models, respectively. It can be observed that by learning to instruct images, the LIT model demonstrates more precise and comprehensive image understanding, as well as stronger robustness against misleading information in the language instructions.
@@ 3.1 Setup
Model Architectures. (1) TinyLLaVA (Zhou et al., 2024): We follow TinyLLaVA to utilize small-scale LLMs, including Qwen-2-0.5B (Yang et al., 2024) and Phi-2-3B (Gunasekar et al., 2023). The SigLIP-400M (Zhai et al., 2023) is adopted as the vision encoder architecture due to its effective performance in combination with small-scale LLMs. The multimodal connector follows LLaVA 1.5 (Liu et al., 2024a) with a two-layer MLP and GELU activation. (2) LLaVA 1.5, LLaVA-NeXT (Liu et al., 2024b): For LLaVA 1.5, we use Vicuna-v1.5-7B (Peng et al., 2023) and Vicuna-v1.5-13B (Peng et al., 2023) as the base LLMs, while for LLaVA-NeXT, Vicuna-v1.5-7B serves as the base LLM. The base vision encoder is CLIP-ViT-L-14 (Radford et al., 2021), and the multimodal connector is a two-layer MLP.
@@
Training Datasets. (1) Pretraining Stage: We adopt the LLaVA 1.5 framework, utilizing the LLaVA-pretrain-558k data (Liu et al., 2024a) for all pretraining phases across TinyLLaVA, LLaVA 1.5, and LLaVA-NeXT. (2) Finetuning Stage: For instruction tuning, we use the LLaVA-mix-665k data (Liu et al., 2024a) on TinyLLaVA and LLaVA 1.5. For LLaVA-NeXT, we use the LLaVA-NeXT-Data (Liu et al., 2024b), an expansion of LLaVA-mix-665k with diverse instruction data. More details can be referred to Section B.2.
@@
Implementation Details. We train all models on NVIDIA A100 GPUs, strictly following the training recipes of TinyLLaVA, LLaVA 1.5 and LLaVA-NeXT. Further details on the hyperparameters can be referred to Section B.1.
@@
Evaluation Benchmarks. We conduct a thorough evaluation of four distinct capabilities using 16 multimodal instruction datasets. These capabilities are categorized as follows: (1) General Visual Question Answering, assessed through VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), ScienceQA (Lu et al., 2022), and VizWiz (Gurari et al., 2018); (2) Comprehensive Multimodal Benchmarks, evaluated using MME (Fu et al., 2023), MMMU (Yue et al., 2024), and MMStar (Chen et al., 2024b); (3) Chart, Document, and OCR Understanding, evaluated using ChartQA (Masry et al., 2022), TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021), and OCR Bench (Liu et al., 2023b); and (4) Image Captioning, assessed through COCO2017 (Lin et al., 2014), Flickr30k (Young et al., 2014), NoCaps (Agrawal et al., 2019), RefCOCO (Kazemzadeh et al., 2014), and TextCaps (Sidorov et al., 2020). The performance metrics used include CIDEr (Vedantam et al., 2015) for all captioning tasks, the perception score for MME, and accuracy for the remaining tasks. The experiments are conducted using LMMs-Eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval/ (Zhang et al., 2024). In  Section 3.4, some tasks are evaluated on the lite version, denoted as “-L”.

### v2

@@ Abstract
We propose L2T, an advancement of visual instruction tuning (VIT). While VIT equips Multimodal LLMs (MLLMs) with promising multimodal capabilities, the current design choices for VIT often result in overfitting and shortcut learning, potentially degrading performance. This gap arises from an overemphasis on instruction-following abilities, while neglecting the proactive understanding of visual information. Inspired by this, L2T adopts a simple yet effective approach by incorporating the loss function into both the instruction and response sequences. It seamlessly expands the training data, and regularizes the MLLMs from overly relying on language priors. Based on this merit, L2T achieves a significant relative improvement of up to 9% on comprehensive multimodal benchmarks, requiring no additional training data and incurring negligible computational overhead. Surprisingly, L2T attains exceptional fundamental visual capabilities, yielding up to an 18% improvement in captioning performance, while simultaneously alleviating hallucination in MLLMs. Github code: https://github.com/Feng-Hong/L2T.
footnotetext: Equal contribution. Work done during Feng Hong’s internship at Microsoft Research Asia.
@@ 1 Introduction
Figure 1: Performance comparison on a broad range of 16 tasks between L2T and VIT using different models, including TinyLLaVA Qwen2-0.5B (Zhou et al., 2024), LLaVA-1.5 Vicuna-7B (Liu et al., 2024a), and LLaVA-1.5 Vicuna-13B (Liu et al., 2024a). The pretraining phase uses the LLaVA-pretrain-558k dataset, while the fine-tuning phase employs the LLaVA-mix-665k dataset.
@@
Typically, visual instruction tuning consists of two stages: pre-training and fine-tuning (Liu et al., 2023a; Zhu et al., 2024). In the pre-training stage, a connector is trained to align visual and language data. The fine-tuning phase is similar to instruction tuning in language models, where the model is trained to generate responses based on multimodal instructions. However, recent research has shown that instruction tuning can lead to a series of issues, such as the knowledge degradation (Ghosh et al., 2024) and hallucinations (Rawte et al., 2023; Tonmoy et al., 2024). Similar problems have also been observed in models that employ visual instruction tuning (Huang et al., 2023; Leng et al., 2024). The potential causes may lie in overfitting and shortcut learning during instruction tuning (Sun et al., 2024a). Figure˜2 illustrates an example of a shortcut, where the model might ignore visual content and generate responses based solely on language priors. Current advancements (Liu et al., 2024a, b; Bai et al., 2023b; Tong et al., 2024) in MLLMs implicitly mitigate such issues by leveraging larger, higher-quality, and more diverse training datasets, larger models, and improved pretrained initializations for visual and language backbones.
In this paper, we propose an orthogonal solution to improving visual instruction tuning: learning to
@@
instruct images as a regularizer, which we refer to as Learning to InstrucT (L2T). Specifically, in addition to learning to generate responses to given images and instructions as usual, L2T also learn to generate instructions for images that exclude templates, which refer to special tokens and high-frequency, low-information template tokens in the instructions. L2T enhances VIT by: (1) expanding the training content to mitigate overfitting without explicitly enlarging the training set (Section˜3.4); and (2) learning to instruct images, which forces the model to focus more on the visual content and prevents learning some shortcuts (Section˜2).
We conduct extensive experiments across a range of 16 tasks, comparing L2T and VIT on different models. As shown in Figure˜1, L2T achieves a significant relative improvement of up to 6% over VIT in overall multimodal task performance. Moreover, L2T shows significant potential in alleviating hallucination issues in MLLMs, as demonstrated across four diverse hallucination benchmarks (Section˜3.3). Through comprehensive and detailed experiments, L2T demonstrates the following advantages: (1) It improves the ability across different multimodal tasks compared to the conventional visual instruction tuning, especially those focusing more on visual content, such as OCR, captioning and hallucination mitigation. (2) Our method is orthogonal to existing research advancements and can further enhance MLLMs’ performance through simple integration. (3) It shows advantages in improving performance while making minimal compromises in training and inference costs.
@@ 2 Method
Problem Formulation. Taking LLaVA as an example, a typical model architecture comprises a pre-trained large language model $f(\cdot)$ with parameters $\theta_{f}$ with the corresponding tokenizer and embedding layer $t(\cdot)$ (e.g., Vicuna (Peng et al., 2023)), a pre-trained visual feature extractor $g(\cdot)$ with parameters $\theta_{g}$ (e.g., CLIP-ViT-L/14 (Radford et al., 2021)), and a cross-modal connector $h(\cdot)$ with parameters $\theta_{h}$, such as a linear layer or MLP. Let $\theta=\{\theta_{f},\theta_{g},\theta_{h}\}$ denote the set of all parameters. For an image $\mathbf{X}_{V}$ and a related text instruction $\mathbf{X}_{I}$, we obtain the corresponding textual response through the following forward process. As shown in Figure˜4, the image $\mathbf{X}_{V}$ is forwarded through the visual feature extractor $g$ and cross-modal connector $h$, mapping it to visual tokens $\mathbf{H}_{V}=h(g(\mathbf{X}_{V}))$ in the language embedding space. $\mathbf{H}_{V}$ is then combined with the language tokens $\mathbf{H}_{I}=t(\mathbf{X}_{I})$ to form a sequence, which is forwarded into LLM $f$ for autoregressive generation of the response sequence $\mathbf{X}_{A}$. The goal of visual instruction tuning is to train the model to exhibit strong multi-modal instruction-following capabilities.
Figure 3: The model architecture using LLaVA as an example, and the data flow for generating responses from images and instructions.
Figure 4: Illustration of L2T. In addition to learning to generate responses like VIT, L2T also learns to generate instructions that exclude templates.
@@
$\mathcal{L}=-\log p_{\theta}(\mathbf{X}_{A}|\mathbf{X}_{V},\mathbf{X}_{I})=-\sum_{i=1}^{L_{A}}\log p_{\theta}(\mathbf{X}_{A,i}|\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A,<i}).$ (1)
@@
Learning to Instruct (L2T). Building upon VIT, we propose extending its paradigm through learning to instruct images. Specifically, in addition to learning how to generate a response given an image and an instruction, we also learn how to generate a meaningful instruction for a given image. For a training sample triplet $(\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A})$, we define the loss function $\mathcal{L}$ as the negative log-likelihood of both the instruction and the response conditioned on the image:
$\mathcal{L}=-\log p_{\theta}(\mathbf{X}_{I},\mathbf{X}_{A}|\mathbf{X}_{V})=\underbrace{-\sum_{i=1}^{L_{I}}\log p_{\theta}(\mathbf{X}_{I,i}|\mathbf{X}_{V},\mathbf{X}_{I,<i})}_{\text{Learn to Instruct}}\underbrace{-\sum_{i=1}^{L_{A}}\log p_{\theta}(\mathbf{X}_{A,i}|\mathbf{X}_{V},\mathbf{X}_{I},\mathbf{X}_{A,<i})}_{\text{Learn to Respond}}.$ (2)
By learning to generate appropriate instructions for images, L2T achieves two key benefits. 1)It naturally expands the data the model learns to fit, helping to alleviate potential overfitting to some extent. 2) Learning to instruct images ensures that the model focuses on the image content, effectively preventing it from ignoring the visual input and relying solely on languages to generate responses.
Template Removal. To ensure the model learns meaningful content related to the image, we exclude the learning of certain irrelevant parts of instructions. Such irrelevant context primarily arises from two sources: (1) system templates and (2) task templates. Specifically, system templates refer to the tokens used to guide the MLLMs in adopting the role of a helpful and polite AI assistant, or as the conversational clues distinguishing whether the content is generated by the “USER” or the “ASSISTANT”. System templates can be easily removed when it is added. Task templates refer to tokens that indicate the task type and output format. We identify task templates by calculating the frequency of all sentences in the entire training dataset and selecting the most frequent ones. The detailed removed task templates can be referred to Appendix˜A. Note that the training data in the pretraining stage is constructed using image-caption pairs, with a set of template prompts serving as instructions, which are unrelated to the image content. Therefore, we apply our method L2T only during the end-to-end fine-tuning stage. In Figure˜4, we present an illustration of L2T.
@@
Figure 5: The visual contribution ($\mathrm{VC}$) distributions of VIT and L2T on the VQAv2 training data and DocVQA test data. Experiments are based on TinyLLaVA Qwen2-0.5B.
(a) Attention of VIT.
(b) Attention of L2T.
Figure 6: Visualization of attention weights. Darker colors indicate higher attention weights. Experiments are based on TinyLLaVA Qwen2-0.5B.
Figure 7: Cases, each including an image, an instruction, and the responses generated by the VIT and L2T models, respectively. The models are based on LLaVA-1.5 Vicuna-7B. (Left) The VIT model shows better OCR capabilities. (Middle) The VIT model demonstrates robustness and effectively avoids being influenced by misleading information in the language instruction. (Right) The VIT model provides a more comprehensive and accurate image description.
Analysis. To assess the rationality of L2T in alleviating shortcut learning in MLLMs, we conduct experiments to gain deeper insight into which aspects of visual or instructional content are most influential in response prediction. Specifically, we first introduce visual contribution ($\mathrm{VC}$), which quantifies the difference in the log-likelihood of the response when conditioned on both the image and instruction, versus when conditioned solely on the instruction:
@@
where $\mathbf{X}_{V}=\emptyset$ in the implementation denotes that the original visual input is replaced with random noise. In this spirit, $\mathrm{VC}$ captures the relative importance of visual signals in response prediction, isolating the contribution of visual information from any confounding effects due to the instructional signals. Figure˜6 presents the VC of L2T and VIT for both the training and testing datasets. Our findings reveal that L2T achieves a substantial relative improvement, with a 9% increase in VC, highlighting its effectiveness in regularizing MLLMs to better utilize visual inputs.
To further substantiate these findings, we visualize the attention weights in L2T and VIT. As shown in Figure˜6, L2T leads to stronger activation in the visual components of MLLM attention heads, indicating more effective utilization of visual inputs. This is consistent with the VC results. More experimental details and additional visualization results can be referred to Appendix˜F.
Moreover, Figure˜7 presents several cases, each consisting of an image, an instruction, and the responses generated by the VIT and L2T models, respectively. It can be observed that by learning to instruct images, the L2T model demonstrates more precise and comprehensive image understanding, as well as stronger robustness against misleading information in the language instructions.
@@ 3.1 Setup
Model Architectures. (1) TinyLLaVA (Zhou et al., 2024): We follow TinyLLaVA to utilize small-scale LLMs, including Qwen-2-0.5B (Yang et al., 2024) and Phi-2-3B (Gunasekar et al., 2023). The SigLIP-400M (Zhai et al., 2023) is adopted as the vision encoder due to its effective performance in combination with small-scale LLMs. The multimodal connector follows LLaVA 1.5 Liu et al. (2024a) with a two-layer MLP and GELU activation. (2) LLaVA 1.5, LLaVA-NeXT (Liu et al., 2024b): For LLaVA 1.5, we use Vicuna-v1.5-7B (Peng et al., 2023) and Vicuna-v1.5-13B (Peng et al., 2023) as the base LLMs, while for LLaVA-NeXT, Vicuna-v1.5-7B serves as the base LLM. The base vision encoder is CLIP-ViT-L-14 (Radford et al., 2021), and the connector is a two-layer MLP.
@@
Training Datasets. (1) Pretraining Stage: We adopt the LLaVA 1.5 framework, utilizing the LLaVA-pretrain-558k data (Liu et al., 2024a) for all pretraining phases. (2) Finetuning Stage: For instruction tuning, we use the LLaVA-mix-665k data (Liu et al., 2024a) on TinyLLaVA and LLaVA 1.5. For LLaVA-NeXT, we use the LLaVA-NeXT-Data (Liu et al., 2024b), an expansion of LLaVA-mix-665k with diverse instruction data. More details can be referred to Section˜B.2.
@@
Implementation Details. We train all models on NVIDIA A100 GPUs, strictly following the training recipes of TinyLLaVA, LLaVA 1.5 and LLaVA-NeXT. See Section˜B.1 for more details.
@@
Evaluation Benchmarks. We conduct a thorough evaluation of four distinct capabilities using 16 multimodal instruction datasets. These capabilities are categorized as follows: (1) General Visual Question Answering, assessed through VQAv2 (Goyal et al., 2017), GQA (Hudson and Manning, 2019), ScienceQA (Lu et al., 2022), and VizWiz (Gurari et al., 2018); (2) Comprehensive Multimodal Benchmarks, evaluated using MME (Fu et al., 2023), MMMU (Yue et al., 2024), and MMStar (Chen et al., 2024b); (3) Chart, Document, and OCR Understanding, evaluated using ChartQA (Masry et al., 2022), TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2021), and OCR Bench (Liu et al., 2023b); and (4) Image Captioning, assessed through COCO2017 (Lin et al., 2014), Flickr30k (Young et al., 2014), NoCaps (Agrawal et al., 2019), RefCOCO (Kazemzadeh et al., 2014), and TextCaps (Sidorov et al., 2020). The performance metrics used include CIDEr (Vedantam et al., 2015) for all captioning tasks, the perception score for MME, and accuracy for the remaining tasks. The experiments are conducted using LMMs-Eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval/ (Zhang et al., 2024). In  Section˜3.4, some tasks are evaluated on the lite version, denoted as “-L".
3.2 Main Results
In Table˜1, we present a comprehensive evaluation of L2T across multiple benchmarks.